# Formation Control via Rotation Symmetry Constraints

### v1

@@ III-A Symmetry-based Control Law
By Proposition 1, note that $V_{0}$ is the vertical stack of the $2\times 2$ blocks $S_{i}\begin{bmatrix}e_{1}&e_{2}\end{bmatrix}=S_{i}$. Since $S_{i}\in SO(2)$ are orthogonal matrices we have
$V_{0}^{T}V_{0}=\sum_{i=1}^{n}S_{i}^{T}S_{i}=nI_{2},$
which shows that the columns of $V_{0}$ are orthogonal. Since $Q$ is PSD and the columns of $V_{0}$ are orthogonal, we can define $\hat{V}$ as an orthonormal eigenbasis $\hat{V}=\begin{bmatrix}\hat{V}_{0}&\hat{V}_{+}\end{bmatrix}$ with $\hat{V}_{0}=\frac{1}{\sqrt{n}}V_{0}$ and $V_{+}$ the orthogonal complement of $V_{0}$ (in other words, $V_{+}$ corresponding to the non-zero eigenvalues $\Lambda_{+}>0$ of $Q$). Hence, $Q$ is equivalent to
$Q=\hat{V}\begin{bmatrix}0&0\\ 0&\Lambda_{+}\end{bmatrix}\hat{V}^{T},$
and the closed-form solution of (7) results in
$p(t)=e^{-Qt}p(0)=\hat{V}\begin{bmatrix}I_{2}&0\\ 0&e^{-\Lambda_{+}t}\end{bmatrix}\hat{V}^{T}p(0).$
Since all non-zero eigenvalues of Q-Q are in OLHP, the dynamics of $p(t)$ exponentially converge to
${\ \lim_{t\to\infty}p(t)=\hat{V}_{0}\hat{V}_{0}^{T}p(0)=\frac{1}{n}V_{0}V_{0}^{\top}p(0)\ }$ (12)
Note that $V_{0}^{T}p(0)=\sum_{i=k}^{n}S_{k}^{T}p_{k}(0)$. Then, the block expression for each agent:
$\lim_{t\rightarrow\infty}p_{i}(t)=\frac{1}{n}S_{i}\sum_{k=1}^{n}S_{k}^{T}p_{k}(0).$
Moreover, observe that $E(\Gamma_{r})^{T}p=0$ implies $p_{i}=S_{i}q$ for some $q\in\mathbb{R}^{2}$. Hence, $\mathrm{Null}(Q)=\mathrm{IM}(V_{0})=\mathcal{F}$, rendering $\mathcal{F}$ exponentially stable as claimed. ∎∎
@@
Consider a MAS consisting of $n=6$ agents, tasked with attaining a $C_{6}$-symmetric configuration (see Fig. 2(c)). Fig. 4 illustrates the underlying $\Gamma$-symmetric graph, with the dashed edge being removed for the chosen communication topology graph $\mathcal{G}_{I}$ with $5$ edges. Note that by using a distance [3] or bearing [5] approach we would require $9$ edges in total to ensure the correct formation shape in $\mathbb{R}^{2}$.
Figure 4: Underlying graph $\mathcal{G}$.
Fig. 5 illustrates the trajectories and resulting configuration obtained by implementing the proposed control law (7), where the corresponding matrix $Q$ is constructed using the rotation elements $\tau(\gamma_{uv})=R(\pi/3)$.
Figure 5: Trajectories generated from (7).
@@ IV Formation Maneuvering
Our main focus has been on achieving and maintaining a target formation shape. Observe in Fig. 5 that due to symmetry the control (7) will successfully drive the agents to a desired formation shape but with respect to an inertial (global) origin. This may be limiting in many practical scenarios where the formation requires the ability to maneuver, that is, to translate, rotate, and scale while preserving the desired shape. To improve the flexibility of a symmetry-based formation control approach, [6] proposed augmenting the closed-loop dynamics for each agent (6) with a virtual state $r(t)$ that enables the agents to agree on a different origin. We leverage this idea to address formation maneuvering as well.
@@
Assume each agent in the MAS has access to a virtual trajectory, predefined by
@@
    a translation $r(t)\in\mathbb{R}^{2}$ with $\dot{r}(t)=v(t)$;
@@
    a rotation $\mathcal{R}(t)\in SO(2)$ with $\dot{\mathcal{R}}(t)=\Omega(t)\mathcal{R}(t)$ where
    $\Omega(t)=\begin{bmatrix}0&-\omega(t)\\ \omega(t)&0\end{bmatrix}$,
    and $\omega(t)$ is the desired angular velocity of the formation;
@@
    a scale factor $s(t)\in\mathbb{R}^{+}$, with $\dot{s}(t)=\alpha(t)s(t)$, $\alpha(t)\in\mathbb{R}$.
Reference trajectories are known a priori in many applications [14]. Therefore, building on Assumption 1 we can define a shifted state for each agent, $c_{i}(t)=p_{i}(t)-r(t)$. Moreover, since the formation is specified with respect to the origin, we define the centroid of the formation at the origin and an axis of rotation passing through it. Under Assumption 1, we then propose the augmented control law
@@
for the agents to converge to the desired configuration while maneuvering along the predefined trajectory.
@@
Consider a MAS consisting of $n$ integrator agents (2) satisfying Assumption 1, whose interaction topology is defined by a spanning tree graph $\mathcal{G}_{I}$, and let
@@
Define $\zeta(t)\in\mathbb{R}^{2n}$ to be the configuration $p(t)\in\mathbb{R}^{2n}$ expressed in a frame moving along the virtual trajectory,
@@
We examine the derivative of each agent $\zeta_{i}(t)\in\mathbb{R}^{2}$ (product rule),
@@
By applying the control (13), we have
$\dot{\zeta}_{i}(t)$ $=-\alpha(t)\zeta_{i}(t)-\Omega(t)\zeta_{i}(t)-\frac{1}{s(t)}\mathcal{R}(t)^{T}v(t)$
$+\frac{1}{s(t)}\mathcal{R}^{T}\big(\sum_{{ij\in\mathcal{E}_{I}}}(\tau({\gamma_{ji}})c_{j}(t)-c_{i}(t))$
$+v(t)+\Omega(t)c_{i}(t)+\alpha(t)c_{i}(t)\big).$
Since $\zeta_{i}(t)=\frac{1}{s(t)}\mathcal{R}(t)c_{i}(t)$, all the trajectory dependent terms cancel, simplifying the expression to
$\dot{\zeta}_{i}(t)=\sum_{{ij\in\mathcal{E}_{I}}}(\tau({\gamma_{ji}})\zeta_{j}(t)-\zeta_{i}(t)).$
This reduces to the analysis of the agents $\dot{\zeta}(t)=-Q\zeta(t)$. By Theorem 1, the dynamics of $\zeta(t)$ ensure that the formation exponentially converges to the set
$\mathcal{F}_{\zeta}=\{\zeta\in\mathbb{R}^{2n}|\tau(\gamma)\zeta_{i}=\zeta_{\gamma(i)},\;\;\forall\gamma\in\Gamma_{r},\;i\in\mathcal{V}\}$
From the definition of $\zeta_{i}(t)$ (14), this set is equivalent to $\mathcal{F}_{c}$, rendering the set $\mathcal{F}_{c}$ exponentially stable as claimed. ∎∎
@@
Consider the same setup as in Example 3 under Assumption 1. A trajectory is predefined to enable the formation to maneuver through obstacles along a desired path. The blue line in Fig. 6 illustrates the trajectory state along the path, and the scaled arrows as the rotation state and scaling with respect to the initial state.
Figure 6: Predefined reference trajectory implemented in control law (13).
Fig. 7 illustrates the resulting trajectories for each agent along the virtual trajectory predefined to each agent under control law (13).
Figure 7: Trajectories generated from (13) along the predefined trajectory.
To further evaluate the system, Fig. 8 illustrates the inter-agent rotation symmetry errors during maneuverings (3). The errors exponentially converge to zero, which shows the effectiveness of the proposed method.
Figure 8: Norm of the symmetry errors
@@ V Extension to $\mathbb{R}^{3}$
The results of Section IV are stated explicitly for formations in $\mathbb{R}^{2}$. Extending these ideas to $\mathbb{R}^{3}$ is relatively straightforward. In this section we provide a numerical example to illustrate how this might be done, with a formal analysis reserved for future work.
Consider a MAS consisting of $n=8$ agents required to achieve a cube formation with an underlying graph $\mathcal{G}$ shown in Fig. 9(a). The target configuration can again be defined as a $\tau(\Gamma)$-symmetric framework, where the rotation elements $\tau(\Gamma)$ are now given by rotation matrices $R\in SO(3)$. In $\mathbb{R}^{3}$, such rotations are defined about coordinate axes, or general hyperplanes through the origin. For a cube formation, a natural choice of symmetries is given by $\mathcal{C}_{4}$ rotations about the coordinate axes (see Fig. 9(b)). For instance, the agents $\{p_{1},\,p_{2},\,p_{3},\,p_{4}\}$ and $\{p_{5},\,p_{6},\,p_{7},\,p_{8}\}$ may each satisfy a $\mathcal{C}_{4}$- symmetric framework about the $z$ axis, and agents $\{p_{2},\,p_{1},\,p_{5},\,p_{6}\}$ can be constrained to form a $\mathcal{C}_{4}$- symmetric framework orthogonal to $z$. These symmetry relations suffice for agents to exchange information according to a communication topology subgraph $\mathcal{G}_{I}$, obtained by removing the dashed edges in Fig. 9. By construction, $\mathcal{G}_{I}$ is a spanning tree. Similar to the planar case, we now define the symmetry-constraining matrix $Q_{z}$ defined by $\mathcal{C}_{4}$ symmetries about the $z$ axis, and $Q_{\perp}$ as the corresponding matrix defined by $\mathcal{C}_{4}$ symmetries about an orthogonal axis of $z$. The resulting symmetry-constraining matrix $Q$ for the cube formation is then obtained as a composition
@@
Figure 9: (a) The underlying graph $\mathcal{G}$, and (b) is the desired $\tau(\Gamma)$-symmetric framework of the cube formation.
By construction, $Q\succeq 0$, and its null-space corresponds to the set of cube-symmetric configurations. Hence the control (7) drives the system exponentially to the desired configuration.
Similar to the planar case, the method can be augmented with a virtual reference state $(r(t),R(t),s(t))$ to achieve maneuvering along a predefined trajectory. Fig. 10 shows the resulted trajectories by implementing control law (13). The agents converge to a cube formation along the predefined trajectory.
Figure 10: Trajectories of the cube formation generated from (13) along a predefined trajectory.
@@ VI Concluding Remarks
We have introduced a formation control method to achieve a target formation by using only rotation symmetry constraints between neighboring agents. Notably, we have demonstrated that a communication spanning tree subgraph with (n1)(n-1) constraints (matching the minimal connectivity requirement) suffices for the implementation of the protocol. By augmenting the control law with a time-varying virtual state we also proved the flexibility of the approach by enabling the ensemble to undergo translations, rotations, and scalings. In addition, we presented an extension of the method to achieve a formation in $\mathbb{R}^{3}$. This work demonstrates the potential of symmetry constraints in addressing formation control problems. Future directions include formally extending the framework to point group elements in $\mathbb{R}^{3}$, exploring directed, and switching graphs, and incorporating leader-follower configurations to generalize the approach to broader classes of configurations and enable fully distributed agreement on time-varying virtual trajectories.

### v2

@@ III-A Symmetry-based Control Law
By Proposition 1, note that $V_{0}$ is the vertical stack of the $2\times 2$ blocks $S_{i}\begin{bmatrix}e_{1}&e_{2}\end{bmatrix}=S_{i}$. Since $S_{i}\in SO(2)$ are orthogonal matrices and $V_{0}^{T}V_{0}=\Sigma_{i=1}^{n}S_{i}^{T}S_{i}=nI_{2}$, then the columns of $V_{0}$ are orthogonal as well. Since $Q$ is PSD and the columns of $V_{0}$ are orthogonal, we can define $\hat{V}$ as an orthonormal eigenbasis $\hat{V}=\begin{bmatrix}\hat{V}_{0}&\hat{V}_{+}\end{bmatrix}$ with $\hat{V}_{0}=\frac{1}{\sqrt{n}}V_{0}$ and $V_{+}$ the orthogonal complement of $V_{0}$. Hence, $Q$ is equivalent to $Q=\hat{V}\left[\begin{smallmatrix}0&0\\ 0&\Lambda_{+}\end{smallmatrix}\right]\hat{V}^{T},$ and the closed-form solution of (7) results in $p(t)=e^{-Qt}p(0)=\hat{V}\left[\begin{smallmatrix}I_{2}&0\\ 0&e^{-\Lambda_{+}t}\end{smallmatrix}\right]\hat{V}^{T}p(0)$. Since all non-zero eigenvalues of Q-Q are in OLHP, the dynamics of $p(t)$ exponentially converge to $p(\infty)=\hat{V}_{0}\hat{V}_{0}^{T}p(0)=\frac{1}{n}V_{0}V_{0}^{\top}p(0)$. Note that $V_{0}^{T}p(0)=\sum_{k=1}^{n}S_{k}^{T}p_{k}(0)$. Then, the block expression for each agent is $p_{i}(\infty)=\frac{1}{n}S_{i}\sum_{k=1}^{n}S_{k}^{T}p_{k}(0).$ Moreover, observe that $E(\Gamma_{r})^{T}p=0$ implies $p_{i}=S_{i}q$ for some $q\in\mathbb{R}^{2}$. Hence, $\mathrm{Null}(Q)=\mathrm{IM}(V_{0})=\mathcal{F}$, rendering $\mathcal{F}$ exponentially stable as claimed. ∎
@@
Consider a MAS consisting of $n=6$ agents, tasked with attaining a $\mathcal{C}_{6}$-symmetric configuration (see Fig. 2(c)). Fig. 3(a) illustrates the underlying $\Gamma$-symmetric graph, with the dashed edge being removed for the chosen communication topology graph $\mathcal{G}_{I}$ with $5$ edges. Note that by using a distance [3] or bearing [5] approach we would require $9$ edges in total to ensure the correct formation shape in $\mathbb{R}^{2}$.
(a) Underlying cycle graph $C_{6}$.
(b) Trajectories generated from (7).
Figure 3: Graph and simulation results for Example 3.
Fig. 3(b) illustrates the trajectories and resulting configuration obtained by implementing the proposed control law (7), where the corresponding matrix $Q$ is constructed using the rotation matrices $\tau(\gamma_{ij})=R(\pi/3)$.
@@ IV Formation Maneuvering
Our main focus has been on achieving and maintaining a target formation shape. Observe in Fig. 3(b) that, due to symmetry, the control (7) successfully drives the agents to a desired formation shape, but with respect to an inertial (global) origin. This may be limiting in many practical scenarios where the formation requires the ability to maneuver, that is, translate, rotate, and scale while preserving the desired shape. To improve the flexibility of a symmetry-based formation control approach, [6] proposed augmenting the closed-loop dynamics for each agent (6) with a virtual state $r(t)$ that enables the agents to agree on a different origin. We leverage this idea to address formation maneuvering as well.
@@
Each agent in the MAS has access to a virtual trajectory predefined by a time-varying
@@
    translation $r(t)\in\mathbb{R}^{2}$ with $\dot{r}(t)=v(t)$;
@@
    rotation $\mathcal{R}(t)\in SO(2)$ with $\dot{\mathcal{R}}(t)=\Omega(t)\mathcal{R}(t)$ where $\Omega(t)=\left[\begin{smallmatrix}0&-\omega(t)\\ \omega(t)&0\end{smallmatrix}\right]$, and $\omega(t)$ is the desired angular velocity of the formation;
@@
    scale factor $s(t)\in\mathbb{R}^{+}$, with $\dot{s}(t)=\alpha(t)s(t)$, $\alpha(t)\in\mathbb{R}$.
Reference trajectories are known a priori in many applications [14]. Therefore, building on Assumption 1, we define a shifted state for each agent, $c_{i}(t)=p_{i}(t)-r(t)$. Moreover, since the formation is specified with respect to a fixed inertial point, we define the centroid of the formation at the origin with the axis of rotation passing through it. Under Assumption 1, we then propose the augmented control law
@@
which enables the agents to converge to the desired configuration while maneuvering along the predefined trajectory.
@@
Consider a MAS consisting of $n$ integrator agents (2) satisfying Assumption 1, whose interaction topology is defined by $\mathcal{G}_{I}$, and let
@@
Define $\zeta(t)\in\mathbb{R}^{2n}$ as the configuration $p(t)\in\mathbb{R}^{2n}$ expressed in a frame moving along the virtual trajectory,
@@
We examine the derivative of each agent state $\zeta_{i}(t)\in\mathbb{R}^{2}$ (product rule),
@@
By applying the control law (11), we have
$\dot{\zeta}_{i}(t)=-\alpha(t)\zeta_{i}(t)-\Omega(t)\zeta_{i}(t)-\frac{1}{s(t)}\mathcal{R}(t)^{T}v(t)+\frac{1}{s(t)}\mathcal{R}^{T}$
$\big(\sum_{{ij\in\mathcal{E}_{I}}}(\tau({\gamma_{ji}})c_{j}(t)-c_{i}(t))+v(t)+\Omega(t)c_{i}(t)+\alpha(t)c_{i}(t)\big).$
Since $\zeta_{i}(t)=\frac{1}{s(t)}\mathcal{R}(t)^{T}c_{i}(t)$, all trajectory dependent terms cancel, simplifying the expression to
$\dot{\zeta}_{i}(t)=\sum_{{ij\in\mathcal{E}_{I}}}(\tau({\gamma_{ji}})\zeta_{j}(t)-\zeta_{i}(t)),$
reducing the analysis of the dynamics to $\dot{\zeta}(t)=-Q\zeta(t)$. By Theorem 1, the dynamics of $\zeta(t)$ ensure that the formation exponentially converges to the set
$\mathcal{F}_{\zeta}=\{\zeta\in\mathbb{R}^{2n}|\tau(\gamma)\zeta_{i}=\zeta_{\gamma(i)},\;\;\forall\gamma\in\Gamma_{r},\;i\in\mathcal{V}\}.$
From the definition of $\zeta_{i}(t)$ (12), this set is equivalent to $\mathcal{F}_{c}$, rendering the set $\mathcal{F}_{c}$ exponentially stable as claimed. ∎
@@
Consider the same setup as in Example 3 under Assumption 1. A trajectory is predefined to enable the formation to maneuver through obstacles along a desired path. The blue line in Fig. 4 illustrates the translational trajectory along the path, and the scaled arrows indicate the rotation and scaling states with respect to the initial state.
Figure 4: Predefined reference trajectory used in control law (11).
Fig. 5 illustrates the resulting agent trajectories along the predefined virtual trajectory under control law (11).
Figure 5: Agent trajectories generated by (11) along the predefined reference trajectory.
To further evaluate the system, Fig. 6 illustrates the inter-agent rotation symmetry errors during maneuvering (3). The errors exponentially converge to zero, showing the effectiveness of the proposed method.
Figure 6: Norm of the symmetry error over time.
@@ V Extension to $\mathbb{R}^{3}$
The results of Section IV are stated explicitly for formations in $\mathbb{R}^{2}$. Extending these ideas to $\mathbb{R}^{3}$ is relatively straightforward. In this section, we provide a numerical example to illustrate how this might be done, with a formal analysis reserved for future work.
Consider a MAS consisting of $n=8$ agents required to achieve a cube formation with an underlying graph $\mathcal{G}$ shown in Fig. 7(a). The target configuration can again be defined as a $\tau(\Gamma)$-symmetric framework, where the elements of $\tau(\Gamma)$ are given by rotation matrices $R\in SO(3)$. In $\mathbb{R}^{3}$, such rotations are defined about coordinate axes, or general hyperplanes through the origin. For a cube formation, a natural choice of symmetries is given by $\mathcal{C}_{4}$ rotations about the coordinate axes (see Fig. 7(b)). For instance, the agents $\{p_{1},\,p_{2},\,p_{3},\,p_{4}\}$ and $\{p_{5},\,p_{6},\,p_{7},\,p_{8}\}$ may each satisfy a $\mathcal{C}_{4}$-symmetric framework about the $z$-axis, and agents $\{p_{2},\,p_{1},\,p_{5},\,p_{6}\}$ can be constrained to form a $\mathcal{C}_{4}$-symmetric framework orthogonal to the $z$-axis. These symmetry relations suffice for agents to exchange information according to the subgraph $\mathcal{G}_{I}$, obtained by removing the dashed edges as shown in Fig. 7. By construction, $\mathcal{G}_{I}$ is also a spanning tree. Similar to the planar case, we now define the symmetry-constraining matrix $Q_{z}$ defined by $\mathcal{C}_{4}$ symmetries about the $z$-axis, and $Q_{\perp}$ as the corresponding matrix defined by $\mathcal{C}_{4}$ symmetries about an axis orthogonal to the $z$-axis. The resulting symmetry-constraining matrix $Q$ for the cube formation is then obtained as a composition
@@
Figure 7: (a) Shows the underlying graph $\mathcal{G}$, and (b) shows the desired $\tau(\Gamma)$-symmetric framework of the cube formation.
By construction, $Q\succeq 0$, and its null-space corresponds to the set of cube-symmetric configurations. Hence the control (7) drives the system exponentially to the desired configuration. Similar to the planar case, this method can be augmented with a virtual reference trajectory $(r(t),\mathcal{R}(t),s(t))$. However, note that in $\mathbb{R}^{3}$, $\Omega(t)$ and $\mathcal{R}(t)$ do not always commute. Hence, during rotations, we also require the matrix $Q$ to undergo a similarity transformation according to $\mathcal{R}(t)$. Fig. 8 shows the resulting trajectories obtained by implementing control law (11). The agents converge to a cube formation along the predefined trajectory.
Figure 8: Cube-formation trajectories generated by (11) along a predefined reference trajectory.
@@ VI Concluding Remarks
This paper introduced a formation control strategy that achieves cyclic target configurations using only rotation symmetry constraints between neighboring agents. In particular, we showed that a communication spanning-tree subgraph with n1n-1 constraints, matching the minimal connectivity requirement, is sufficient to implement the protocol. By augmenting the control law with a time-varying virtual state, we further demonstrated that the formation can perform coordinated translations, rotations, and scalings. We also presented a numerical extension of the method to formations in $\mathbb{R}^{3}$. Overall, the results highlight the potential of symmetry-based constraints for formation control. Future work will focus on a formal extension to broader point-group elements in $\mathbb{R}^{3}$, directed and switching interaction graphs, and leader-follower architectures, with the goal of enabling fully distributed agreement on time-varying virtual trajectories